# Deep Learning based Key Information Extraction from Business Documents: Systematic Literature Review

### v1

@@ D. Overview
  • Input modalities: Indicates, which modalities derived from input documents were considered for KIE. Besides textual, visual and layout-oriented features, we also include custom hand-crafted features in the analysis.
@@
  • Independent of external OCR: Indicates whether the KIE approach requires no external OCR engine. If this is the case, the system is either responsible both for text reading and information extraction (end-to-end) or alternatively requires no text reading stage at all and maps input documents immediately to desired outputs (OCR-free) (Sassioui et al., 2023).
@@
12 grid 461 Invoices, receipts 1,05
@@
[…]
@@
24 sequence 1540 Receipts, NDAs $\sim similar-to\sim$ 536
@@
[…]
47 sequence 2467 Reciepts, exams n/a
@@
78 graph 199 Forms 6,2
79 sequence 1739 Forms, receipts, NDAs 362
80 graph 973 Receipts 0,0536
81 sequence 2712 Forms, Receipts 175000
@@
89 sequence 1199 Forms, receipts 794
@@ E. Sequence-based approaches
43 word 1105000 RVL-CDIP*, UCSF Industry Documents Library, Common Crawl* Masked Visual-Language Modeling T5 U-Net n/a With relative attention biases
@@
52 word 43000000 IIT-CDIP*, Custom Unidirectional LM, Bidirectional LM, Sequence-to-Sequence LM LayoutXLM n/a Custom
@@
86 word 6000000 IIT-CDIP Masked Detection Modeling BERT ResNet-50, Deformable DETR Custom Vision-Language Decoder Deformable cross-attention & language-conditioned cross-attention
@@ F. Graph-based approaches
22 Learned soft adjacency matrix sentence Transformer-embeddings + ResNet50-embeddings 1024 undirected Horizontal and vertical distance + aspect ratio + relative height + relative width + sentence length ratio GCN on node-edge-node-triplets with MLP 1 BiLSTM-CRF
@@
26 Iteratively learned via NN text lines CNN-encodings + normalized coordinates + detection confidence + class prediction 256 directed CNN-embeddings + normalized coordinates + class prediction + distance GCN + multi-head attention 7;7;4 Node classification
@@
85 Custom layout tree; nearest neighbors in 4 major directions text lines / n/a directed / Custom graph reordering & graph masking 1 Sequence labeling layer
@@
94 Fully connected word / n/a directed / Token paths predicted (binary classification) 1 Greedy search
@@ G. Grid-based approaches
40 word Image height / 8 Image width / 8 256 Word embeddings (BERT/ RoBERTa/ LayoutLM) + ResNet18-FPN-embeddings all-zero vector Binary classifier followed by set of binary classifiers, Auxiliary Semantic Segmentation
48 token vertical range of bboxes vertical range of bboxes 256 ResNet18-embeddings + horizontal and vertical relative coordinates all-zero vector BiLSTM
@@
H. Evaluation
Table ST7. Analysis of evaluation methodologies
@@ 20 F1 Custom
21 Precision, Recall, F1 FUNSD, SROIE 0,7927 0,9524
@@
23 n/a n/a F1 SROIE, Custom 0,9618
24 n/a n/a Precision, Recall, F1 CORD, FUNSD, Kleister-NDA 0,9699 0,8455
25 n/a n/a Accuracy, Precision, Recall, F1 SROIE, Custom 0,8920
26 Precision, Recall, F1 FUNSD 0,6652
27 F1 CORD, Kleister-NDA, Kleister-Charity, SROIE, SROIE* 0,9441 0,9817
28 F1 CORD, FUNSD 0,9894 0,8796
@@ 31 F1, nTED, A/B test Custom
32 F1 CORD, CORD*, FUNSD, Custom 0,9250 0,7160
@@ 35 Precision, Recall, F1 Custom
36 F1 FUNSD, Custom 0,5722
37 n/a n/a Precision, Recall, F1 FUNSD 0,8514
38 n/a n/a F1 FUNSD 0,8366
39 n/a n/a Precision, Recall, F1 EPHOIE, FUNSD, SROIE 0,8309 0,9688
40 F1 SROIE, Custom 0,9640
41 n/a n/a F1 CORD, Custom 0,9571
42 n/a n/a Accuracy, Precision, Recall, F1 CORD, FUNSD, SROIE 0,9477 0,8109 0,9303
43 F1 CORD, SROIE 0,9633 0,9810
@@ 44 Accuracy, F1 MARG, NIST, Tobacco, Custom
45 n/a n/a F1 EPHOIE, FUNSD, SROIE 0,8133 0,9657
@@ 46 F1 Custom
47 n/a n/a F1 EPHOIE, SROIE 0,9612
48 n/a n/a F1 EPHOIE, SROIE, Custom 0,9654
49 F1 CORD, FUNSD, Kleister-NDA, SROIE 0,9601 0,8420 0,9781
50 n/a n/a Precision, Recall, F1 FUNSD, Custom 0,6596
@@ 51 F1 Custom
52 F1 CORD, EPHOIE, FUNSD*, LastDoc4000, SROIE, XFUND* 0,9684 0,9790
53 Precision, Recall, F1 CORD, FUNSD*, SROIE, SROIE* 0,9745 0,9821
@@
55 n/a n/a Precision, Recall, F1 FUNSD 0,8652
56 n/a n/a F1 FUNSD, XFUND 0,8335
57 n/a n/a Precision, Recall, F1 CORD, FUNSD, SROIE* 0,9728 0,8452
58 n/a n/a F1 CORD, FUNSD 0,9746 0,9208
59 TED-Accuracy, F1 CORD, Ticket, Custom 0,8410
60 n/a n/a F1 CORD, FUNSD, Custom 0,9728 0,8469
61 n/a n/a F1 FUNSD, CORD 0,9700 0,4610
@@ 59 TED-Accuracy, F1 CORD, Ticket, Custom 0,8410
63 n/a n/a F1 CORD, FUNSD, SROIE, Kleister-NDA 0,9721 0,9312 0,9755
64 n/a n/a Precision, Recall, F1 CORD, FUNSD, EPHOIE, XFUND 0,9607 0,8415
65 n/a n/a F1 FUNSD, CORD, SROIE 0,9738 0,8649 0,9791
@@ 66 Tag Error Rate Custom
68 n/a n/a Precision, Recall, F1 FUNSD, Custom 0,8346
69 n/a n/a F1 SROIE 0,9794
70 n/a n/a F1 SROIE 0,9822
71 Precision, Recall, F1 CORD, FUNSD, SROIE 0,9575 0,8345 0,9740
72 GAnTED, F1 FUNSD, NAF, IAM NER 0,6500
73 n/a n/a F1 CORD, WildReceipt 0,9541
74 Precision, Recall, F1 EPHOIE, SCID, SROIE 0,9693
75 F1, Document Accuracy Rate CORD, SROIE, Custom 0,8450 0,8730
@@ 76 F1 VAIPE
77 n/a n/a F1 CORD, WildReceipt 0,9665
78 F1 FUNSD 0,8225
79 n/a n/a F1 FUNSD, CORD, Kleister-NDA 0,9783 0,9101
80 Precision, Recall, F1 SROIE 0,9363
81 F1 CORD, FUNSD, SROIE 0,9412 0,9032 0,9788
82 F1 FUNSD, XFUND 0,8079
83 n/a n/a F1 SROIE, POIE 0,8587
84 n/a n/a Precision, Recall, F1 CORD, FUNSD, SROIE, Custom 0,9737 0,8635 0,9831
85 n/a n/a Precision, Recall, F1 CORD, FUNSD, XFUND 0,9775 0,9439
86 F1 CORD, FUNSD 0,9820 0,8400
87 n/a n/a F1 CORD, FUNSD 0,$97$ 0,9286
88 n/a n/a F1 CORD, FUNSD 0,9749 0,9376
89 F1 CORD, DeepForm, FUNSD, Kleister-Charity, PWC 0,9758 0,9162
90 n/a n/a F1 CORD, FUNSD, SROIE 0,9719 0,9320 0,9727
91 n/a n/a Precision, Recall, F1 CORD, FUNSD 0,9693 0,8877
92 n/a n/a F1 CORD, FUNSD, SIBR, XFUND 0,9565 0,9112
@@ 94 F1 CORD-r, FUNSD-r
95 F1 FUNSD, SROIE 0,7527 0,9610
96 n/a n/a F1 CORD, FUNSD, SROIE 0,9693 0,8795 0,9845

### v2

@@ D. Overview
  • Input modalities: Indicates, which modalities derived from input documents were considered for KIE. Besides textual, visual and layout-oriented features, we also include custom hand-crafted features in the analysis. Note that in case of LLM-based approaches, we do not consider the textual prompts to the LLM (required by design) as a text-based modality.
@@
  • Independent of external OCR: Indicates whether the KIE approach requires no external OCR engine.
@@
12 grid 461 Invoices, receipts 1.05
@@
16 15310 Invoices, receipts $\sim$ n/a
17 sequence 3500 Banking documents n/a $\sim$
18 sequence 219476 Purchase orders $\sim$1,4
@@
24 sequence 1739 Receipts, NDAs $\sim$ $\sim$ 536
@@ 44 $\sim$1010630 Forms, journals, reports n/a
46 $\sim$15000 Banking documents, invoices n/a
47 sequence 2467 Receipts, exams n/a
@@ 46 $\sim$15000 Banking documents, invoices n/a
78 graph 199 Forms 6.2
79 graph 973 Receipts 0.0536
80 LLM 2712 Forms, Receipts 175000
@@
88 sequence 1199 Forms, receipts, reports 794
@@ 92 LLM 3800 Forms, reports 7000
97 sequence 2712 Forms, receipts, NDAs 425
98 sequence 2567 Forms, invoices, personnel archives n/a
99 graph 199 Forms 14
100 graph + sequence 2592 Receipts, forms n/a
101 LLM 4172 Forms, receipts 7000
102 LLM 1199 Forms, receipts 7368
103 graph 5305 Forms, receipts n/a
104 graph 3912 Forms, receipts 0.8
105 sequence 1739 Forms, receipts, NDAs 362
106 LLM 3800 Forms, reports 8100
107 sequence 5966 Forms n/a
108 graph 2534 Contracts n/a
109 sequence 1172 Forms, receipts 114
110 sequence 2742 Purchase orders n/a
111 sequence 1592 Forms 32
112 1199 Forms, invoices n/a
113 LLM 1199 Receipts, reports 7000
114 LLM 2172 Forms, receipts 7000
115 graph + sequence 4581 Forms, tickets 182
116 LLM 3556 Forms, invoices, receipts n/a
117 LLM 17793 Forms, invoices, receipts n/a
118 LLM 1973 Receipts 7000
119 sequence 1990 Forms 419
120 sequence 3520 Forms, invoices, receipts 14.2
121 graph 1439 Receipts n/a
122 sequence 1199 Forms, receipts 125
123 sequence 6405 Forms, invoices, receipts n/a
124 sequence 1973 Receipts n/a
125 graph + sequence 1592 Forms 160
126 LLM 10582 Forms, receipts, reports $\sim$ $\sim$7853
127 sequence 3666 Forms, receipts, exams n/a
128 graph + sequence 1199 Forms, receipts 154
129 sequence 4224 Receipts 21
130 sequence 3226 Forms, receipts 131
@@ E. Sequence-based approaches
43 word 1105000 RVL-CDIP*, UCSF Industry Documents Library, Common Crawl* Masked Visual-Language Modeling T5 U-Net T5 With relative attention biases and contextualized image embeddings
@@
52 word 43000000 IIT-CDIP*, Custom Unidirectional LM, Bidirectional LM, Sequence-to-Sequence LM LayoutXLM n/a Transformer
@@
85 word 6000000 IIT-CDIP Masked Detection Modeling BERT ResNet50, Deformable DETR Custom Vision-Language Decoder Deformable cross-attention & language-conditioned cross-attention
@@
97 text line 1000000 RVL-CDIP, FUNSD, CORD, SROIE, Kleister-NDA, DocVQA, PubLayNet Masked Language Modeling, Document Categorization, Layout Segment Categorization, Reading Order Identification of Layout Segments, Relation Extraction RoBERTa / Sequence labeling layer
98 word + segment / / / LayoutXLM ResNet50 MLP
100 word 6000000 IIT-CDIP Masked Visual-Language Modeling, Key Point Location, Cross-modal Alignment Identification RoBERTa, InfoXLM / / disentangled attention mechanism
107 word / / / LayoutXLM ResNet18 Custom Relation Prediction
109 word / / / LayoutLM / Sequence labeling layer & binary classifier for hierarchical elements
110 word / / / LayoutLMv2 / k-NN search with prototypes
111 word / / / LiLT[InfoXLM] / Efficient Global Pointer bi-directional attention complementation mechanism
115 word / / / Transformer ResNet50 Sequence labeling layer multimodal self-attention mechanism
119 word / / / LayoutLMv3 LayoutLMv3 Score Pointer SCA attention
120 word / / / GRU + Transformer / Sequence labeling layer
122 word 50000 Industry Document Library OCR Token Prediction, Bounding Box Prediction Transformer Transformer Sequence labeling layer
123 word 9624 SVRD, DocILE Token classification, token order prediction LayoutLMv3, LayoutMask LayoutLMv3, LayoutMask Token classification + token order prediction
124 / n/a Curved SynthText, ICDAR 2013,ICDAR 2015, MLT 2017, Total-Text, TextOCR, HierText, COCO Text, Open Image Spatial-Window Prompting, Prefix-Window Prompting / Swin Transformer + FPN Transformer decoder
125 word / / / LiLT[EN-R], LiLT[InfoXLM] ResNet101 Token classification + token link
127 word 6000000 IIT-CDIP Masked Visual-Language Modeling, Key Point Location, Relative Position Classification, Word-Patch Alignment RoBERTa, InfoXLM Transformer Sequence labeling layer dual-stream attention enhancement mechanism
128 word + segment 10000 IIT-CDIP Masked Language Modeling LayoutLMv3 / Graph node classification
129 segment 18000 Custom Question-Answering Linear embedding layers, GRU ResNet50 Sequence labeling layer Layout-Self-Attention
130 token / / / LiLT[RoBERTa], LiLT[InfoXLM], LiLT[XLMRoBERTa] / SPADE
@@ F. Graph-based approaches
22 Learned soft adjacency matrix sentence Transformer-embeddings + ResNet50-embeddings 1024 undirected Horizontal and vertical distance + aspect ratio + relative height + relative width + sentence length ratio GCN on node-edge-node-triplets with MLP and self-attention 1 BiLSTM-CRF
@@
26 Iteratively learned via NN text lines CNN-embeddings + normalized coordinates + detection confidence + class prediction 256 directed CNN-embeddings + normalized coordinates + class prediction + distance GCN + multi-head attention 7;7;4 Node classification
@@
84 Custom layout tree; nearest neighbors in 4 major directions text lines / / directed / Custom graph reordering & graph masking 1 Sequence labeling layer
@@
94 Fully connected word / / directed / Token paths predicted (binary classification) 1 Greedy search
@@
99 k-NN segment Coordinates + area of coordinates + regional encoding 24&1465 undirected Angle + euclidean distance + relative coordinates + U-net-embeddings GNN & GAT 2&2 Node classification
100 k-NN (k=100 & 50) word + segment + region Coordinates + height + width 192 n/a horizontal and vertical distance GAT 12 n/a
103 Links nodes if there are no other nodes in between them word BERT-embeddings with mapping layer n/a directed Angle GCN n/a Node classification
104 k-NN (k=4) neighbors with heighest weight word Word2Vec-embeddings + CNN-embeddings + coordinates + euclidean distance n/a undirected Normalized Euclidean distance + normalized Chebyshev distance + normalized Manhattan distance + normalized mutual information between image segments + correlation coefficient between image segments + cosine text similarity GCN n/a Node classification
108 k-NN (k=6) & Max-distance connection segment BiLSTM-embeddings n/a directed Absolute position ratio + aspect ratio GCN on node-edge-node-triplets with MLP and self-attention n/a Node classification
115 Fully connected within same category word / / undirected / / n/a /
121 Nearest neighbors in 4 major directions segment n/a n/a undirected n/a ChebConv-GCN 12 Node classification
125 Fully connected within text box token / / directed / / / Token classification + token link
128 Fully connected & learned relationship matrix segment BERT-embeddings + normalized coordinates + center coordinates + width + height n/a directed / GCN 1 Node classification
@@ G. Grid-based approaches
40 word Image height / 8 Image width / 8 256 Word embeddings (BERT/RoBERTa/LayoutLM)* + ResNet18-FPN-embeddings all-zero vector Binary classifier followed by set of binary classifiers, Auxiliary Semantic Segmentation
48 token maximal - minimal y coordinate of normalized bounding boxes maximal - minimal x coordinate of normalized bounding boxes 256 ResNet18-embeddings + horizontal and vertical relative coordinates all-zero vector BiLSTM
@@
H. LLM-based approaches
Table ST7. Analysis of LLM-based approaches
# Backbone Image encoding Pre-Training KIE Fine-Tuning
LLM(s) Open source/weights Encoder Resolution # instructions Datasets Tasks # instructions Datasets Tasks PEFT # trained parameters (M)
80 Davinci-003, ChatGPT / / / / / / / / /
92 mPLUG-Owl mPLUG-Owl 224x224 (20 crops) / / / n/a DeepForm, Kleister-Charity Unified downstream task (VQA, IE, NLI, Image Captioning), Text Reading, Key Points Generation LoRA 86
101 Vicuna-7B Swin Transformer 2560×2560 1520000 Common Crawl*, Custom Text detection, text recognition, text spotting, paragraph reading, full-text reading, image captioning 370000 FUNSD, SORIE, POIE Text detection, text recognition, text spotting, paragraph reading, full-text reading, image captioning 7000
102 Llama-7B LayoutLMv3 n/a / / / n/a CORD, FUNSD Summarization, Document QA 7000
106 mPLUG-Owl2 ViT/L-14 448x448 (9 crops) 4036402 DocStruct4M Document/Webpage Parsing, Table Parsing, Chart Parsing, Natural Image Parsing, Multi-grained Text Localization 56809 DeepForm, Kleister-Charity Document Parsing 7000
113 LLaMA-2-7B Swin Transformer 1536x1536 / / / n/a CORD, SROIE, DeepForm, Kleister-Charity document classification, information extraction, visual question answering, optical character recognition, visual grounding, image captioning, table reconstruction LoRA n/a
114 Vicuna-7B-v1.5 LayoutLMv3 n/a 5700000 PubLayNet, DocLayNet, DocBank, RVL-CDIP, DocILE Document Dense Description, Text and Layout Reconstruction, Document Layout Analysis, Table Understanding, Masked vision language modeling, position masking, geometric pre-training 300000 DocILE Information extraction 7000
116 CodeLlama / / / / / n/a CORD, VRDU Information extraction LoRA n/a
117 PaLM 2-S, Gemini Pro / / / / / n/a Custom, CORD, Payment, VRDU Information extraction n/a
118 LLaMA2-7B-Chat LayoutLMv3 n/a / / / 25877 DocReason25k VQA 7000
126 Falcon-1B, Llama2-7B / / n/a IIT-CDIP, DocBank Block infilling 274845 BuDDIE, CORD, DeepForm, FUNSD, Kleister-Charity, PWC, SROIE, VRDU* Information extraction 7000
I. Evaluation
Table ST8. Analysis of evaluation methodologies
@@ 20 F1 Custom
21 Precision, Recall, F1 FUNSD, SROIE 0.7927 0.9524
@@
23 n/a n/a F1 SROIE, Custom 0.9618
24 n/a n/a Precision, Recall, F1 CORD, FUNSD, Kleister-NDA 0.9699 0.8455
25 n/a n/a Accuracy, Precision, Recall, F1 SROIE, Custom 0.8920
26 Precision, Recall, F1 FUNSD 0.6652
27 F1 CORD, Kleister-NDA, Kleister-Charity, SROIE, SROIE* 0.9441 0.9817
28 F1 CORD, FUNSD 0.9894 0.8796
@@ 31 F1, nTED, A/B test Custom
32 F1 CORD, CORD*, FUNSD, Custom 0.9250 0.7160
@@ 35 Precision, Recall, F1 Custom
36 F1 FUNSD, Custom 0.5722
37 n/a n/a Precision, Recall, F1 FUNSD 0.8514
38 n/a n/a F1 FUNSD 0.8366
39 n/a n/a Precision, Recall, F1 EPHOIE, FUNSD, SROIE 0.8309 0.9688
40 F1 SROIE, Custom 0.9640
41 n/a n/a F1 CORD, Custom 0.9571
42 n/a n/a Accuracy, Precision, Recall, F1 CORD, FUNSD, SROIE 0.9477 0.8109 0.9303
43 F1 CORD, SROIE 0.9633 0.9810
@@ 44 Accuracy, F1 MARG, NIST, Tobacco, Custom
45 n/a n/a F1 EPHOIE, FUNSD, SROIE 0.8133 0.9657
@@ 46 F1 Custom
47 n/a n/a F1 EPHOIE, SROIE 0.9612
48 F1 EPHOIE, SROIE, Custom 0.9654
49 F1 CORD, FUNSD, Kleister-NDA, SROIE 0.9601 0.8420 0.9781
50 n/a n/a Precision, Recall, F1 FUNSD, Custom 0.6596
@@ 51 F1 Custom
52 F1 CORD, EPHOIE, FUNSD*, LastDoc4000, SROIE, XFUND* 0.9684 0.9790
53 Precision, Recall, F1 CORD, FUNSD*, SROIE, SROIE* 0.9745 0.9821
@@
55 n/a n/a Precision, Recall, F1 FUNSD 0.8652
56 n/a n/a F1 FUNSD, XFUND 0.8335
57 n/a n/a Precision, Recall, F1 CORD, FUNSD, SROIE* 0.9728 0.8452
58 n/a n/a F1 CORD, FUNSD 0.9746 0.9208
59 TED-Accuracy, F1 CORD, Ticket, Custom 0.8410
60 n/a n/a F1 CORD, FUNSD, Custom 0.9728 0.8469
61 n/a n/a F1 FUNSD, CORD 0.9700 0.4610
@@
63 n/a n/a F1 CORD, FUNSD, SROIE, Kleister-NDA 0.9721 0.9312 0.9755
64 n/a n/a Precision, Recall, F1 CORD, EPHOIE, FUNSD, XFUND 0.9607 0.8415
65 n/a n/a F1 FUNSD, CORD, SROIE 0.9738 0.8649 0.9791
@@ 66 Tag Error Rate Custom
68 n/a n/a Precision, Recall, F1 FUNSD, Custom 0.8346
69 n/a n/a F1 SROIE, Custom 0.9822
70 n/a n/a F1 SROIE 0.9794
71 Precision, Recall, F1 CORD, FUNSD, SROIE 0.9575 0.8345 0.9740
72 GAnTED, F1 FUNSD, NAF, IAM NER 0.6500
73 n/a n/a F1 CORD, WildReceipt 0.9541
74 Precision, Recall, F1 EPHOIE, SCID, SROIE 0.9693
75 F1, Document Accuracy Rate CORD, SROIE, Custom 0.8450 0.8730
@@ 76 F1 VAIPE
77 n/a n/a F1 CORD, WildReceipt 0.9665
78 F1 FUNSD 0.8225
79 Precision, Recall, F1 SROIE 0.9363
80 F1 CORD, FUNSD, SROIE 0.9412 0.9032 0.9788
81 F1 FUNSD, XFUND 0.8079
82 n/a n/a F1 SROIE, POIE 0.8587
83 n/a n/a Precision, Recall, F1 CORD, FUNSD, SROIE, Custom 0.9737 0.8635 0.9831
84 n/a n/a Precision, Recall, F1 CORD, FUNSD, XFUND 0.9775 0.9439
85 F1 CORD, FUNSD 0.9820 0.8400
86 n/a n/a F1 CORD, FUNSD 0.9797 0.9286
87 n/a n/a F1 CORD, FUNSD 0.9749 0.9376
88 F1 CORD, DeepForm, FUNSD, Kleister-Charity, PWC 0.9758 0.9162
89 n/a n/a F1 CORD, FUNSD, SROIE 0.9719 0.9320 0.9727
90 n/a n/a Precision, Recall, F1 CORD, FUNSD 0.9693 0.8877
91 n/a n/a F1 CORD, FUNSD, SIBR, XFUND 0.9565 0.9112
92 F1 DeepForm, Kleister-Charity
@@ 94 F1 CORD-r, FUNSD-r
95 F1 FUNSD, SROIE 0.7527 0.9610
96 n/a n/a F1 CORD, FUNSD, SROIE 0.9693 0.8795 0.9845
97 n/a n/a F1 CORD, FUNSD, SROIE, Kleister-NDA 0.9804 0.9306 0.9794
98 n/a n/a Precision, Recall, F1 Custom, SIBR, XFUND
99 n/a n/a F1 FUNSD 0.6476
100 n/a n/a Precision, Recall, F1 CORD, FUNSD, XFUND 0.9638 0.8852
101 Accuracy FUNSD, SROIE, POIE
102 F1 CORD, FUNSD 0.9860 0.9530
103 n/a n/a F1 CORD, FUNSD, WildReceipt, SROIE, XFUND 0.9660 0.8150 0.9220
104 n/a n/a F1 CORD, FUNSD, WildReceipt, SROIE 0.9720 0.8040 0.9070
105 n/a n/a F1 CORD, FUNSD, Kleister-NDA 0.9783 0.9101
106 F1 DeepForm, Kleister-Charity
107 F1 Custom, SIBR
108 n/a n/a F1 Custom
109 n/a n/a F1 CORD, FUNSD, SROIE 0.8730 0.8320
110 n/a n/a Precision, Recall, F1 Custom
111 n/a n/a F1 FUNSD, XFUND 0.9011
112 F1 RFUND, SIBR
113 F1 CORD, Kleister-Charity, SROIE 0.8930 0.9100
114 ANLS CORD, FUNSD, SROIE
115 n/a n/a Precision, Recall, F1 CER-VIR, FUNSD, XFUND 0.8728
116 F1, nTED CORD, VRDU
117 F1 CORD, VRDU 0.9557
118 n/a n/a Accuracy CORD, SROIE
119 n/a n/a Precision, Recall, F1 Custom, FUNSD, XFUND 0.9274
120 n/a n/a F1 Custom
121 Accuracy, Precision, Recall, F1 Custom
122 n/a n/a Precision, Recall, F1 CORD, FUNSD 0.9662 0.9141
123 n/a n/a F1 CORD-r, FUNSD-r, EPHOIE, SIBR, SROIE, WildReceipts, XFUND 0.9761
124 TED-Accuracy, F1 CORD, SROIE 0.8480 0.8560
125 n/a n/a F1 FUNSD, XFUND 0.7202
126 F1 BuDDIE, CORD, DeepForm, FUNSD, Kleister-Charity, PWC, SROIE, VRDU* 0.6740 0.5180 0.9190
127 n/a n/a Precision, Recall, F1 CORD, EPHOIE, FUNSD, SROIE, XFUND 0.9734 0.9265 0.9872
128 n/a n/a F1 CORD, FUNSD 0.9108 0.9650
129 n/a n/a F1 CORD, Custom, SROIE, XFUND 0.9674 0.9751
130 n/a n/a F1 CORD, Custom, FUNSD 0.9639 0.7899